SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (Feb.7, 2006)

Attendees: Rainer Bartoldus, Kris Chaisanguanthum, Al Eisner, Fang Fang, Arik Kreisel, Selina Li, Rahmat Rahmat, Dong Su, Dan Walker

Operations

  • PEP's vacumm leak problem is still not fixed even after the long access last week. BaBar is still taking physics data at a lumi of ~4x10^33. The trigger system didn't really have any problems for the past week.

EMT

  • We are still not sure about the error bars in the EMT deadtime plot obtained from the ambient database.
  • Su Dong wonders if adding up the sub-systems' deadtime would give us the total deadtime. Of course there is a correlation between some systems. Ideally, we should look at the VMON data by crate. The ambient database uses Epics data, not VMON. Rainer while on shift noticed that there is still some visible amount of deadtime due to EMT (even we have been using the prescale factor of 3 right now). Maybe it's not just the FEX that contributed to the deadtime. Perhaps we should get a table of FEX time for each system from Chris O'Grady.

DCT/DCZ

  • A DQM shifter noticed that the ZPD tandip distribution for SL7 and SL10 looked different from the reference plots. See plots. Selina went back a few runs and found that it looked like the distributions changed shape at the beginning of February. The ZPD is sensitive to background, and in particular, it looked like we seem to have more LER background now. This perhaps relates to the LER vacumm leak? Selina is currently studying the L1 accepts to see if we can help identify where the leak is. Kris checked and confirmed that the references we used for L1T JAS plots are from July data. Rainer pointed out that the beam current is more balanced between HER and LER now than before, and perhaps this contributed to the more symmetric distribution of the tandip.
  • The alarming messages printed out in the ROM logfile were greatly reduced since last Wednesday. See a summary here.
  • Marc might come to SLAC in late Feb or early March to work on the TSF firmware.

L3

  • Rainer has fixed a bug due to a limitation on the number of TSF segments in the L3 tracking code, see HN posted. Note that this is independent of Arik's L3 tracking studies.
  • The L3TFastMonPostProc for producing post-run plots was installed by Jim Hamilton on Feb.6, 2006 Monday day shift. With this implementation, the run by run and montly plots are available right after each run is taken. The next item in discussion is to create a web page to put these plots for easy access and then decide whether or not to use the L1T fastmon archive mechanism or put these plots in the electronic logbook. There is a proposal to put the monitoring plots in the log book for all sub-systems.
  • Al asked about keeping history of the run info releases for L3. Run 3 is missing and Run 5 is remained to be done. Maybe Prafulla can work on this for Run 3 and Run 5 or should comment on this.

Data Quality and Monitoring

  • In run 5 data quality check, Rahmat noticed that the TRG plots were missing for OPR, and got all of them back afer pointed out to Mark (the DQG manager).
  • Brian Lindquist pointed out that changing "Fast_Recent" to "Medium_Full" will enable the first 45 minutes of a run to be displayed rather than only the first 900 seconds for the L1TFMon time history plots.

This page is created by Selina Li (selina@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.